# CORE: Lossless Compression for Retrieval-Augmented LLMs via Reinforcement Learning

### v1

@@ 3.2 Main Results
The detailed comparison results are presented in Table 1. Each experiment was repeated five times, and the average results are reported. We compared the following approaches: a baseline RAG method without compression using the top 1, 3, 5, and 10 documents prepended to the original input; a traditional retrieval method (BM25) for compression; off-the-shelf LLMs (Qwen2.5-1.5B-Instruct and DeepSeek-V3 (671B)); and state-of-the-art trained compressors for RAG (RECOMP-Abstractive, RECOMP-Extractive, and NoiseFilter-IB). For a fair comparison, RECOMP-Abstractive, RECOMP-Extractive, NoiseFilter-IB, and our method CORE were all trained based on the same model, Qwen2.5-1.5B-Instruct. All of these trained compressors were trained using five documents, and we report their performance both on five-document compression and their generalization to ten-document compression.
@@
As shown in the table, prepending retrieved documents yields a substantial improvement over using no retrieval, indicating that the retrieved documents contain valuable information relevant to the answer. Performance improves as the number of retrieved documents increases, albeit with diminishing returns, which can be attributed to the decreasing relevance of lower-ranked documents. When compressing the top five documents, all baseline compression methods result in performance degradation to varying degrees compared to the uncompressed baseline. Specifically, the traditional BM25 approach leads to a significant drop in performance—for instance, a decrease of 13 EM points on NQ. Using the off-the-shelf Qwen2.5-1.5B-Instruct model for compression improves results over BM25 but still underperforms relative to the uncompressed baseline. Among the trained compressors, both RECOMP and NoiseFilter-IB exhibit a decline of 3–5 EM points compared to the uncompressed setting. We also observe that when using the extremely large-scale DeepSeek-V3 model (with 671B parameters) for compression, near-lossless compression is achieved on TriviaQA and HotpotQA; however, performance on NQ and 2Wiki remains below that of the uncompressed baseline. In contrast, our method—with only 1.5B parameters—achieves lossless compression across all four datasets and even outperforms the uncompressed baseline by 1 to 7 EM points. Overall, our approach not only significantly outperforms compression methods of comparable model size, but also maintains an advantage over the hundreds-of-times-larger DeepSeek model, demonstrating that it achieves high compression ratios without sacrificing performance—and even enhancing it.
@@
Moreover, when generalizing the trained compressors to top-10 documents, the aforementioned conclusions remain valid. Our method continues to perform best. On NQ, it achieves a compression ratio of 3.6% in token usage while improving Exact Match (EM) by 3.2 points compared to prepending all ten documents. Similar trends are observed on TQA, with a compression ratio of 2.3% and a gain of 2.4 EM points relative to using the full set of ten documents. On HotpotQA, the method attains a compression ratio of 2.5% and an improvement of 0.7 EM points. On 2WikiMultihopQA, it achieves a compression ratio of 3.1% along with a notable increase of 6.9 EM points over prepending all ten documents.
3.3 Generalization Ability Analysis
@@
  • Model-Agnostic Compression Framework. Figure 5 presents a comparison of compressors trained using different models across the four datasets, while the baseline LLM used to generate the final answer remains fixed as Qwen2.5-14B-Instruct. The compressor models evaluated include LLama3.2-1B-Instruct, Qwen2.5-1.5B-Instruct, and LLama3.2-3B-Instruct, which differ in both model architecture and number of parameters. The results demonstrate that: (1) Compressors trained with our method using any of these three models achieve lossless compression and outperform the uncompressed baseline, indicating that our framework is not dependent on a specific compressor architecture and attesting to the robustness of our approach. (2) As the size of the compressor model increases, so does the compression performance, reflecting the scaling law.
    Table 2: Transfer results to Llama-3.1-8B-Instruct.
@@
  • Training-Free Adaptation to New LLMs. We report the results of transferring our compressor to the LLaMA-3.1-8B model in Table 2, using exactly the same input (including in-context examples and prepended document summaries) as employed with the Qwen2.5-14B-Instruct model. Note that the compressor was trained using feedback generated by Qwen2.5-14B-Instruct. The results indicate that the trained compressor baselines (e.g., RECOMP) exhibit limited generalization ability, as reflected in their increased performance gap relative to the baseline of prepending full documents. In contrast, our method demonstrates strong generalization, outperforming the baseline of prepending full documents across all four datasets while maintaining a high compression rate. This suggests that the summaries generated by our trained compressor are inherently high-quality and contain key information relevant to the answer, enabling effective transfer to other LLMs to support accurate answering.
3.4 Further Analysis
@@
  • The Impact of Distillation and GRPO. Table 3 presents an ablation study on the two key components of our method: distillation and GRPO. Here, “w/o distillation” denotes training the compressor with GRPO directly, bypassing the warm-start phase, while “w/o GRPO” indicates using only the distillation step without subsequent reinforcement learning. The results demonstrate that removing either component leads to performance degradation, confirming the necessity of both. The decline is more substantial when GRPO is omitted, highlighting the crucial role of reinforcement learning in the absence of explicit supervision. Distillation injects external knowledge into the model, providing a favorable starting point for RL training and thereby enabling RL to more effectively unlock the model’s full potential.
@@
  • The Impact of $\alpha$. Figure 6 illustrates the performance of our method under different values of $\alpha$, where $\alpha$ controls the weighting coefficient of the F1 reward term. Setting $\alpha=0$ corresponds to using only the Exact Match (EM) reward. Results show that performance improves when $\alpha>0$ across all datasets, indicating the effectiveness of the F1 reward in alleviating the sparsity issue of the EM reward. Although the optimal value of $\alpha$ varies across datasets, values between 0.2 and 0.5 generally yield strong performance.
@@
3.5 Case Study
To conduct an in-depth analysis of the advantages of our compressor, we performed case studies on one single-hop QA dataset (NQ) and one multi-hop QA dataset (2Wiki), with the results presented in Table 4 and Table 5, respectively. For each case, we compared the summaries generated by off-the-shelf Qwen2.5-1.5B, RECOMP, and our method CORE based on the same set of documents, as well as the final predictions made by the LLM after prepending these summaries.
As shown in the tables, although the summaries produced by off-the-shelf Qwen2.5-1.5B are concise, they largely fail to capture key information relevant to answering the question. In contrast, RECOMP demonstrates better summarization capability but is prone to being overwhelmed by lengthy documents, resulting in misjudgments and even generating misleading information—such as the statement in Table 4: “The U.S. stopped drafting for the Vietnam War after the Selective Service System was officially abolished in December 1972”—which leads the LLM to produce the incorrect answer “1972”. Our method, CORE, accurately extracts answer-critical information from lengthy documents, exemplified by the summary: “The U.S. stopped drafting for the Vietnam War in 1973 after announcing the decision by Secretary of Defense Melvin Laird earlier that year,” thereby enabling the LLM to generate the correct answer “1973”. This indicates that our compressor, trained with an end-task target-oriented reinforcement learning optimization strategy, can produce document summaries that are most helpful for answering the given question while effectively filtering out irrelevant information.
4 Related Work
4.1 Compression in RAG
RAG enhances the performance of Large Language Models (LLMs) on knowledge-intensive tasks by retrieving the most relevant documents from extensive knowledge bases and prepending them as contextual information to the original input Ram et al. (2023); Fan et al. (2024). However, this approach requires the LLM to process significantly longer token sequences, resulting in increased computational costs. To mitigate this issue, researchers have begun to explore methods for compressing retrieved documents in retrieval-augmented generation systems Xu et al. ; Cao et al. (2024); Jin et al. (2024b); Zhu et al. (2024); Kim & Thorne (2025); Wu et al. (2025); Jin et al. (2024a); Zhang et al. (2024a). Xu et al. propose compressing retrieved documents into a textual summary prior to in-context augmentation, training the compressor via data selection and distillation. Cao et al. (2024) introduce a Query-Guided Compressor (QGC), which leverages queries to guide the context compression process, thereby effectively preserving key information. Jin et al. (2024b) refine retrieved documents into Key Supporting Evidence (KSE) through knowledge synthesis, supervised fine-tuning (SFT), and preference alignment. Zhu et al. (2024) propose an information-theoretic approach that filters noise by simultaneously maximizing the mutual information between the compressed content and the ground-truth output. Kim & Thorne (2025) train a compressor to extract critical information using reward functions based on heuristic rules. However, most of these methods compromise end-task performance and lack end-to-end optimization, thus failing to ensure that the compressor maximally benefits the LLM.
4.2 Reinforcement Learning
Reinforcement learning has recently achieved significant success, enabling LLMs to develop reasoning capabilities without explicit step-by-step supervision, relying instead on rule-based reward signals Liu et al. (2024); Shao et al. (2024). Inspired by these advances, several studies have leveraged reinforcement learning (RL) to enhance retrieval-augmented generation (RAG). For instance, Kulkarni et al. (2024) employ RL to autonomously decide whether to retrieve documents or not. Zhang et al. (2024b) use RL to optimize the ranking of retrieved documents. Mao et al. (2024) propose a framework for training query rewriting models without human annotations for RAG. Chen et al. (2025b) improve retrieval-augmented generation through multi-agent reinforcement learning, incorporating a Query Rewriter, Retriever, and Generator. Additionally, a line of research applies RL to integrate search with reasoning in a step-by-step manner. Chen et al. (2025a) introduce a novel framework named ReSearch, which trains LLMs to reason with search via reinforcement learning without relying on supervised data for reasoning steps. Similar approaches include R1-Searcher Song et al. (2025), WebThinker Li et al. (2025), and DeepResearcher Zheng et al. (2025). In contrast to these works, our approach utilizes RL to train a document compressor in an end-to-end fashion, enhancing the efficiency of RAG while preserving its performance.

### v3

@@ 3.2 Overall Performance
The detailed comparison results are presented in Table 1. We evaluate the following approaches: traditional RAG without compression using the top 1, 3, 5, and 10 documents prepended to the original input; BM25-based compression; off-the-shelf LLMs (Qwen2.5-1.5B-Instruct and DeepSeek-V3); and state-of-the-art trainable RAG compressors (RECOMP-Abstractive, RECOMP-Extractive, and NoiseFilter-IB). For fair comparison, all trainable methods were trained using the same backbone model, Qwen2.5-1.5B-Instruct. These compressors were trained on five-document inputs, and we report their performance both on in-domain five-document compression and out-of-domain generalization to ten-document compression.
  • RAG versus No RAG. As shown in Table 1, prepending the original input with retrieved documents yields a substantial improvement over the no-retrieval baseline, indicating that these documents provide valuable information for generating the answer. Performance improves as the number of retrieved documents increases from 1 to 10, albeit with diminishing returns—a trend attributable to the decreasing relevance of lower-ranked documents. However, this performance gain comes at the cost of a significant increase in the number of additional tokens the LLM must encode, rising from 0 to over 1,400 tokens.
  • Lossless Compression of CORE. Our compressor was trained using a context of five documents. As presented in Table 1, compared to prepending the full content of all five documents to the original input (i.e., the uncompressed baseline), our method achieves a high compression rate of approximately 6% with no loss in performance. Remarkably, on all four datasets, the approach not only maintains performance but also enhances EM by 1 to 7 points. For example, on NQ, the model achieves an EM of 41.02 with compressed input, surpassing the uncompressed score of 38.03.
@@
  • Comparison with Compression Baselines. When compressing the top five documents, all baseline compression methods result in performance degradation to varying degrees compared to the uncompressed baseline. Specifically, BM25 leads to a substantial performance drop. Using the off-the-shelf Qwen2.5-1.5B-Instruct model for compression yields better results than BM25 but still underperforms relative to the uncompressed baseline. Surprisingly, even the large-scale DeepSeek-V3 model (with 671B parameters) achieves only near-lossless compression on TriviaQA and HotpotQA, while performance on NQ and 2Wiki remains below the uncompressed baseline. As for the trained compressors, both RECOMP and NoiseFilter-IB exhibit a decline of 3–5 EM points compared to no compression. In contrast, our method, CORE, delivers the best performance—not only significantly outperforming compression methods of comparable model size, but also maintaining an advantage over the hundreds-of-times-larger DeepSeek model, demonstrating the benefit and importance of end-to-end optimization. More detailed results can be found in Tables 6 and 7.
@@
  • Generalization to 10-Doc Compression. When generalizing the trained compressors to handle the top-10 documents without retraining, the aforementioned conclusions remain valid. CORE continues to achieve lossless compression and performs best among all compression methods. On NQ, it achieves a token compression ratio of 3.6% while improving the EM by 3.2 points compared to using all ten documents. Similar trends are observed on TriviaQA, with a compression ratio of 2.3% and a gain of 2.4 EM points relative to the full-document baseline. On HotpotQA, CORE achieves a compression ratio of 2.5% and an improvement of 0.7 EM points. For 2WikiMultihopQA, it obtains a compression ratio of 3.1% along with a notable increase of 6.9 EM points.
3.3 Robustness and Generalization Ability Analysis
@@
  • Robustness of CORE Across Compressor Architectures. To evaluate whether the effectiveness of CORE depends on the choice of compressor model, we compared the performance of compressors trained using different model architectures—while keeping the downstream LLM (Qwen2.5-14B-Instruct) fixed. The compressors tested include LLaMA3.2-1B-Instruct, Qwen2.5-1.5B-Instruct, and LLaMA3.2-3B-Instruct, which vary in architecture and parameter count. As shown in Figure 3, the results indicate that: (1) These trained compressors consistently achieve lossless compression and outperform the uncompressed baseline (represented by the red reference line in the figure, which corresponds to prepending the full document content), confirming that our training framework is robust and not tied to a specific compressor architecture. (2) Compression performance improves as the size of the compressor model increases, consistent with scaling laws.
    Table 2: Zero-shot transfer of the the trained compressor to Llama-3.1-8B-Instruct.
@@
  • Transferability of CORE Across Downstream LLMs. We evaluate the transferability of our trained compressor and other baseline compressors to a new downstream LLM, the LLaMA-3.1-8B model. The results are presented in Table 2. Note that all trainable compressors, including our own, were trained using feedback generated by Qwen2.5-14B-Instruct. The findings reveal that existing trained compressors (e.g., RECOMP) exhibit limited generalization capability, as indicated by a larger performance gap relative to the baseline of prepending full documents. In contrast, CORE demonstrates stronger generalization, achieving lossless compression on the new downstream LLM. Specifically, it not only maintains performance but also surpasses the full-document baseline (i.e., no compression) across all four datasets while retaining a high compression rate. These results suggest that the summaries produced by our method are inherently high-quality and preserve key information necessary for accurate answering, thereby enabling effective transfer to other LLMs.
@@ 3.4 Ablation Study
Table 3 presents an ablation study on the two stages of our method: distillation and GRPO. Here, “w/o distillation” denotes training the compressor with GRPO directly, bypassing the warm-start phase, while “w/o GRPO” indicates using only the distillation step without subsequent reinforcement learning. The results demonstrate that removing either component leads to performance degradation, confirming the necessity of both. In addition, the decline is more substantial when GRPO is omitted, highlighting the crucial role of reinforcement learning in the absence of explicit supervision. Distillation injects external knowledge into the model, providing a favorable starting point for RL training and thereby enabling RL to more effectively unlock the compressor’s full potential.
3.5 The Impact of $\alpha$.
@@
Figure 4 illustrates the performance of our method under different values of $\alpha$, which controls the weighting coefficient of the F1 reward term. Setting $\alpha=0$ corresponds to using only the EM reward. Results indicate that performance improves when $\alpha>0$ across all datasets, demonstrating the effectiveness of the F1 reward in mitigating the sparsity issue associated with the EM reward. However, the optimal value of $\alpha$ is dataset-dependent; values between 0.2 and 0.5 generally yield strong performance.
3.6 Case Study
To conduct an in-depth analysis of the advantages of our compressor, we performed case studies on one single-hop QA dataset (NQ) and one multi-hop QA dataset (2Wiki), with the results presented in Table 4 and Table 5, respectively. For each case, we compared the summaries generated by off-the-shelf Qwen2.5-1.5B-Instruct, RECOMP, and our method CORE based on the same set of documents, as well as the predicted answers generated by the LLM after prepending these summaries. As shown in the tables, although the summaries produced by off-the-shelf Qwen2.5-1.5B are concise, they largely fail to capture key information relevant to answering the question. In contrast, RECOMP demonstrates better summarization capability but is prone to being overwhelmed by lengthy documents, resulting in misjudgments and even generating misleading information—such as the statement in Table 4: “The U.S. stopped drafting for the Vietnam War after the Selective Service System was officially abolished in December 1972”—which leads the downstream LLM to produce the incorrect answer “1972”. Our method, CORE, accurately extracts answer-critical information from lengthy documents, exemplified by the summary: “The U.S. stopped drafting for the Vietnam War in 1973 after announcing the decision by Secretary of Defense Melvin Laird earlier that year”, thereby enabling the LLM to generate the correct answer “1973”. This indicates that our compressor, trained with an end-task target-oriented reinforcement learning optimization strategy, can produce document summaries that are most helpful for answering the given question while effectively filtering out irrelevant information.
4 Related Work
  • Context Compression in RAG. RAG enhances the performance of LLMs on knowledge-intensive tasks by retrieving the most relevant documents from extensive knowledge bases and prepending them as contextual information to the original input (Ram et al., 2023; Fan et al., 2024; Lin et al., 2023; Shi et al., 2023). However, this approach requires the LLM to process significantly longer token sequences, resulting in increased computational costs. To mitigate this issue, researchers have begun to explore methods for compressing retrieved documents in RAG systems (Xu et al., 2024; Cao et al., 2024; Jin et al., 2024b; Zhu et al., 2024; Kim & Thorne, 2025; Rau et al., 2024; Wu et al., 2025; Louis et al., 2025; Jin et al., 2024a; Li et al., 2024a; b; Zhang et al., 2024a). For instance, Xu et al. (2024) propose compressing retrieved documents into textual summaries before in-context augmentation, training the compressor through data selection and distillation. Similarly, Cao et al. (2024) introduce a Query-Guided Compressor (QGC) that uses queries to guide the compression process, effectively preserving essential information. Jin et al. (2024b) refine retrieved documents into Key Supporting Evidence (KSE) through a combination of knowledge synthesis, supervised fine-tuning (SFT), and preference alignment. Meanwhile, Zhu et al. (2024) present an information-theoretic approach called NoiseFilter-IB, which filters noise by maximizing the mutual information between the compressed content and the ground-truth output. Additionally, Kim & Thorne (2025) train a compressor to extract critical information using reward functions based on predefined heuristic rules. However, most of these methods are heuristic in nature, and due to the lack of ideal compression labels, the compressed content they produce cannot be guaranteed to benefit downstream LLMs. In contrast, our method, CORE, adopts an end-to-end optimization approach to address these limitations.
  • Reinforcement Learning. Reinforcement learning (RL) has recently achieved notable success, enabling LLMs to develop reasoning capabilities without explicit step-by-step supervision (Liu et al., 2024; Shao et al., 2024; Guo et al., 2025). Building on these advances, several studies have applied RL to improve RAG (Ke et al., 2024). For example, Kulkarni et al. (2024) use RL to autonomously decide whether to retrieve documents, while Zhang et al. (2024b) employ RL to optimize the ranking of retrieved documents. Similarly, Mao et al. (2024) propose a framework for training query rewriting models for RAG without relying on human annotations. MMOA-RAG (Chen et al., 2025b) enhances RAG through multi-agent reinforcement learning, incorporating a query rewriter, retriever, and generator. Moreover, a line of research has utilized RL to integrate search with reasoning in a step-by-step manner (Singh et al., 2025). For instance, Chen et al. (2025a) introduce a framework called ReSearch, which trains LLMs to reason with search using RL, without requiring supervised data for reasoning steps. Related approaches include R1-Searcher (Song et al., 2025), WebThinker (Li et al., 2025), and DeepResearcher (Zheng et al., 2025). In contrast to these studies, our approach leverages RL to train a document compressor in an end-to-end manner, improving the efficiency of RAG while maintaining high performance.
5 Conclusion
This paper analyzes the limitations of current context compression methods for RAG. A primary challenge is the lack of optimal reference summaries for supervised learning, which often results in performance degradation in downstream tasks. To overcome this, we frame context compression as a reinforcement learning problem, utilizing downstream task performance as a reward signal to train the compression policy, thereby enabling end-to-end optimization. Extensive experiments demonstrate that our proposed method, CORE, achieves effectively lossless compression by maintaining a high compression ratio while preserving original task performance. Surprisingly, CORE not only preserves but actually enhances performance on all downstream tasks. Further in-depth analysis provides additional insights into the efficacy of our approach.
Appendix A Use of LLMs
Large language models (LLMs) were employed solely as writing assistants to enhance the language, improve clarity, and check grammatical correctness. They were not used to generate research ideas, design or implement methodologies, conduct data analysis, or produce any of the results presented in this work. The authors assume full responsibility for the entire content of the paper.
Appendix B Prompt Templates
Figure 5 displays the prompt employed to train the compressor model for generating a summary of the retrieved documents, conditioned on the given query. This prompt is notably concise. For end-task answer generation, the prompt provided to the LLM $M$ is illustrated in Figure 6, which incorporates few-shot in-context examples, the (generated summary of) retrieved documents, and the question.
Figure 5: Prompt template used to generate the summary.
Figure 6: Prompt template for LLM QA.
Appendix C Case Study
To conduct an in-depth analysis of the advantages of our compressor, we performed case studies on one single-hop QA dataset (NQ) and one multi-hop QA dataset (2Wiki), with the results presented in Table 4 and Table 5, respectively. For each case, we compared the summaries generated by off-the-shelf Qwen2.5-1.5B-Instruct, RECOMP, and our method CORE based on the same set of documents, as well as the predicted answers generated by the LLM after prepending these summaries.
@@ Appendix C Case Study
Appendix D Impact of Different Compressors on Performance
In our previous experiments, we employed Qwen2.5-1.5B as the initial model to train our compressor. In this section, we utilize two additional models—Llama3.2-1B and Llama3.2-3B—as starting points to train our compressor and the baseline compressor, respectively. The experimental results are presented in Table 6 and Table 7. As shown in the results, our method CORE continues to achieve lossless compression with both models, maintaining a high token compression ratio while exhibiting no performance degradation in terms of Exact Match (EM) and F1 score compared to uncompressed RAG. Furthermore, under both new model configurations, our approach consistently outperforms the baseline methods, indicating that its superiority is not dependent on a specific model architecture and thus demonstrates strong robustness.
We also observe that our method adheres to a form of scaling law: the compressor trained using the 3B model outperforms the one trained with the 1B model. Specifically, the 1B compressor improves performance by 1–4 EM points over the uncompressed baseline, while the 3B compressor yields gains of 3–9 EM points.
Table 6: Open-domain QA results using Qwen2.5-14B-Instruct as the downstream LLM ($M$). The reported token counts represent the length of in-context documents, excluding few-shot examples. RECOMP and our method CORE are both trained using llama3.2-1B-Instruct.
NQ TriviaQA HotpotQA 2WikiMultihopQA
EM F1 # tok EM F1 # tok EM F1 # tok EM F1 # tok
No Retrieval 0.2136 0.3097 0 0.5323 0.5998 0 0.2105 0.2948 0 0.2611 0.2951 0
RAG without compression
Top1 Document 0.3446 0.4441 142 0.6082 0.6870 143 0.2920 0.3893 147 0.2679 0.3187 153
Top3 Documents 0.3778 0.4845 427 0.6260 0.7102 430 0.3164 0.4187 442 0.2789 0.3358 460
Top5 Documents 0.3803 0.4916 712 0.6410 0.7248 715 0.3299 0.4369 737 0.2964 0.3521 766
Top10 Documents 0.3867 0.5003 1425 0.6440 0.7292 1428 0.3395 0.4488 1471 0.3104 0.3675 1531
Compression of top 5 docs
llama3.2-1B 0.3147 0.4227 64 0.5552 0.6415 60 0.2648 0.3639 58 0.2498 0.3003 61
Deepseek-V3 (671B) 0.3773 0.5039 54 0.6528 0.7433 51 0.3359 0.4483 48 0.2507 0.3031 45
RECOMP (1B) 0.3410 0.4655 57 0.6071 0.6880 48 0.2987 0.4121 49 0.3045 0.3653 33
CORE (1B) 0.3947 0.4923 47 0.6483 0.7287 43 0.3344 0.4454 45 0.3378 0.3969 34
Compression of top 10 docs (with the compressor trained on top 5 docs)
llama3.2-1B 0.3141 0.4228 62 0.5651 0.6512 58 0.2663 0.3661 56 0.2493 0.3006 61
Deepseek-V3 (671B) 0.3779 0.5107 56 0.6529 0.7445 53 0.3462 0.4569 50 0.2900 0.3464 40
RECOMP (1B) 0.3421 0.4661 59 0.6095 0.6917 52 0.2982 0.4105 55 0.3072 0.3681 44
CORE (1B) 0.4033 0.5033 47 0.6521 0.7296 45 0.3412 0.4500 48 0.3586 0.4162 42
Table 7: Open-domain QA results using Qwen2.5-14B-Instruct as the downstream LLM ($M$). The reported token counts represent the length of in-context documents, excluding few-shot examples. RECOMP and our method CORE are both trained using llama3.2-3B-Instruct.
NQ TriviaQA HotpotQA 2WikiMultihopQA
EM F1 # tok EM F1 # tok EM F1 # tok EM F1 # tok
No Retrieval 0.2136 0.3097 0 0.5323 0.5998 0 0.2105 0.2948 0 0.2611 0.2951 0
RAG without compression
Top1 Document 0.3446 0.4441 142 0.6082 0.6870 143 0.2920 0.3893 147 0.2679 0.3187 153
Top3 Documents 0.3778 0.4845 427 0.6260 0.7102 430 0.3164 0.4187 442 0.2789 0.3358 460
Top5 Documents 0.3803 0.4916 712 0.6410 0.7248 715 0.3299 0.4369 737 0.2964 0.3521 766
Top10 Documents 0.3867 0.5003 1425 0.6440 0.7292 1428 0.3395 0.4488 1471 0.3104 0.3675 1531
Compression of top 5 docs
llama3.2-3B 0.3252 0.4334 60 0.5650 0.6521 59 0.2772 0.3809 58 0.2485 0.2995 60
Deepseek-V3 (671B) 0.3773 0.5039 54 0.6528 0.7433 51 0.3359 0.4483 48 0.2507 0.3031 45
RECOMP (3B) 0.3657 0.4912 55 0.6183 0.6920 47 0.3025 0.4238 52 0.3274 0.3806 42
CORE (3B) 0.4310 0.5234 32 0.6650 0.7306 38 0.3507 0.4736 51 0.3905 0.4474 40
Compression of top 10 docs (with the compressor trained on top 5 docs)
llama3.2-3B 0.3318 0.4359 61 0.5720 0.6588 57 0.2791 0.3854 60 0.2491 0.3011 59
Deepseek-V3 (671B) 0.3779 0.5107 56 0.6529 0.7445 53 0.3462 0.4569 50 0.2900 0.3464 40
RECOMP (3B) 0.3682 0.4963 52 0.6205 0.6973 44 0.3077 0.4261 54 0.3312 0.3869 50
CORE (3B) 0.4526 0.5467 33 0.6736 0.7404 37 0.3693 0.4926 51 0.4071 0.4633 48